# Corner functions from entanglement indices of harmonic lattices

### v1

@@ Abstract
We study the entanglement indices such as logarithmic negativities (LNs) and mutual informations (MIs) between two adjacent subsets in a isolated universal set $U$ of harmonic oscillators arranged on a two dimensional lattice within a sufficiently large square. First, we verify the values of the corner functions of angle $\pi/2,\pi/4,3\pi/4$ presented in the previous study which adopts periodic boundary conditions (PBCs) for the $U$. The values of each corner function obtained from LNs are nearly equal to those in the previous ones, while those of $3\pi/4$ from MIs are not sufficiently consistent with those computed from LNs. Next, for the case where the universal system $U$ satisfies the fixed end boundary conditions(FBCs), we calculate LNs, MIs at several locations in $U$, compare them, especially corner functions with the values obtained in the PBCs case, and examine the effect of the fixed ends. In addition, we examine Renyi entropies for sets of three dimensional lattice sites, the corner functions and the edge terms with solid angles and dihedral angles $\pi/2,\pi/4$, respectively.
@@ 1 Introduction
Other quantity that indicates the degree of entanglement between A and its environment B is Renyi entropy[2], which is an extension of the entanglement entropy and is expressed by:
@@
In the case that two subsystems $A_{1}$ and $A_{2}$ are interior of the universal system $U$: $U\supset A=A_{1}+A_{2}$, there are two representative measures of entanglement between $A_{1}$ and $A_{2}$:logarithmic negativity (LN) and mutual information(MI)[3, 4]. The LN is given by
@@
After the seminal work by Hawking and Bekenstein[5, 6], which showed that the entropy of a black hole is proportional to the area of its event horizon, the so-called holographic principle was proposed which asserted that quantum theoretical processes in $d_{U}$ dimensional space could be explained by processes in one smaller dimension,i.e.,the surface area[7, 8]. Furthermore, in the two dimensional space($d_{U}=2$), when the boundary between two adjacent subsystems contains sharp corners of opening angles $\theta_{k},k=1,2,\ldots$, it is said that the many quantities related to the entanglement can be expanded as
$\alpha_{B}P_{B}/\delta+\sum_{k}\beta(\theta_{k})\log P_{B}/\delta+\ldots,$ (5)
@@
where $P_{B}$ is the the boundary length between $A$ and $U$ or between $A_{1}$ and $A_{2}$, $\delta$ is the UV cutoff[9]. The coefficient $\alpha_{B}$ correspondent to the area law depends on the cut-off $\delta$. On the other hand, the corner functions $\beta(\theta_{k})$s are universal, independent of the way the continuous limit to the underlying theory is taken. In three dimensional space $d_{U}=3$, when the boundary surface with area $P_{B}$ has edges with length $L_{j}$ and dihedral angles $\theta_{j},(j=1,2,\ldots)$ as well as sharp vertices with solid angles $\Omega_{k}$, the expansion is given by
$\alpha_{B}P_{B}/\delta^{2}+\sum_{k}\beta(\Omega_{k})\log P_{B}/\delta^{2}+\sum_{j}e(\theta_{j})L_{j}/\delta+\ldots,$ (6)
where edge term coefficients $e(\theta_{j})$ are also universal[10].
@@
For a two dimensional(2D) harmonic lattice with periodic boundary conditions (PBCs), which is a discretized massless scalar field in a torus-like space, the Ref[11] shows that the values of the corner functions of $\theta=\pi/2,\pi/4,3\pi/4$ obtained from LNs are highly consistent with those of conformal field theory describing the scalar field. It is also claimed that the values of the corner functions obtained from MIs corresponding to the Renyi entropies of order 1/2 are consistent with the above LN-derived values. In the computing process, we take the relation between the scale $L$ of the universal set $U$, the scale $l$ of $A$ and the lattice spacing $a$ to be $L\gg l\gg a$, so that the calculation on the lattice can approximate the based field theory which is the continuous limit. With those scale conditions, even when the canonical variables on the lattice sites of $U$ satisfy the fixed-edge boundary conditions(FBCs), the values of those corner function in the region sufficiently far from the boundary of $U$ are expected to be consistent with the above ones calculated under the PBCs.
@@
In this paper, according to Ref.[11], we first consider, as $A_{2}$, a set of lattice sites in several figures with corners of open angles $\theta=\pi/2,\pi/4,3\pi/4$ on their boundaries and scales defined by the number of sites on the shortest side ranging from six to a maximum of about thirty. We suppose that $A$ is a set of sites in a square whose scale is 3 to 4 times larger than that of $A_{2}$ and sites of the set $A_{1}=A-A_{2}$ encloses $A_{2}$ sites. We calculate LNs and MIs between $A_{1}$ and $A_{2}$ for each figure while changing the scales of $A_{2}$ and A. We fit those quantities by the formula (5) to obtain the corner functions and their standard deviations in order to verify those values in the reference. Next, for the case where the canonical variables on the lattice sites satisfy the FBCs, we perform the same procedure as before at many locations in $U$ and observe whether the values of each corner function near the center of $U$ are consistent with those under the PBCs and how they change approaching the boundary of $U$.
@@
This paper is organized as follows: In section 2, we take the Hamiltonian of harmonic lattice which is the discretized massless scalar field. We present analytical solutions $p_{\vec{r}}$ and $q_{\vec{r}}$ of the Hamilton equations under PBCs and FBCs, as well as their vacuum correlation functions $X=\braket{0|q_{\vec{r}},q_{\vec{s}}|0}$ and $P=\braket{0|p_{\vec{r}},p_{\vec{s}}|0}$, where $\vec{r}$ and $\vec{s}$ are position vectors of lattice sites. Given the specific shapes and scales of the universal set $U$ and subsets $A$, $A_{1}$, $A_{2}$, after reviewing the prescription to compute LN and various entropies from $X$ and $P$ matrices, we obtain the value of each corner function of $\theta=\pi/2,\pi/4,3\pi/4$ (and those standard deviation) for harmonic lattice with PBCs and examine whether they are consistent with the previous study. In section 3, we compute the corner functions at many locations in $U$ which is a set of lattice sites under FBCs by the same computational prescription in Section 2, and analyze them. In section 4, we examine, in a 3D lattice, the corner functions and the edge term coefficients, respectively, of solid angles and of dihedral angles $\pi/2$$\pi/4$ of Renyi entropy of order $1/2$. In section 5, we present some conclusions.
@@ 2 Models of harmonic lattices and entanglement indices computation
and correlation functions of generalized coordinate $q_{\vec{r}}$ and their conjugate momentum $p_{\vec{r}}$ expressed as follows[11]:
@@
For the ground state of (21), the values of correlators (14),(15), (19),(20) are also to be obtained from the matrix $K$ as follows [12, 13]:
@@
Consider a set of sites inside the universal system $U$ described in the previous section, and denote it as a subsystem $A$. Also, let its environment B=UAB=U-A. Using the correlators (14),(15), (19),(20), or (23),(24), we can obtain the Renyi entropies $S_{A}^{(n)}$and entanglement entropy S(A)S(A)[14, 15]. First, we make the matrices of rows and columns of the above correlators restricted to those corresponding to the sites of the subsystem, i.e., $X_{\vec{u}\vec{v}}$ and $P_{\vec{u}\vec{v}}$, where $\vec{u},\vec{v}$ are the positions of sites in the subsystem $A$. Second, we compute the matrix $\Xi=\sqrt{XP}$ and its eigenvalues $\xi_{i}$. Using the eigenvalues, we can describe the entanglement entropy(2) and Renyi entropies(3) as follows:
@@
$S_{A}^{(n)}$ $=\frac{1}{1-n}\sum_{i=1}^{\dim(\mathcal{H}_{A})}\log\left((\xi_{i}+\frac{1}{2})^{n}-(\xi_{i}-\frac{1}{2})^{n}\right)$ (26)
Furthermore, we divide the subsystem $A$ into subsubsystems $A_{1}$ and $_{2}, A=A$$A_{1}+A_{2}$. Let us consider mutual information (MI) and logarithmic negativity (LN) as indices of the entanglement between $A_{1}$ and $A_{2}$[4, 11]. The MI and corresponding combination of Renyi entropies are given by
@@
From the matrices $X,P$, we can also calculate LN between the subsystems. First, we find the partial transposition of the density matrix $\rho_{A}$ with respect to $A_{2}$, which corresponds to the time reversal only to the degrees of freedom in $A_{2}$[16]. Therefore, it does not change the matrix $X$. On the otherhand, as for the matrix $P$, when we rewrite the matrix into the form as:
@@
where $\tilde{b}(\theta_{i}^{(j)})$ and $b(\theta_{i}^{(j)})$ are so called corner functions and $\theta^{(j)}_{i}$ is the angle corresponding to the $i$-th vertex of $SB$ belongs to $A_{j}$[11].
@@
We define the scale $l$ and its perimeter $P_{B}$ of each shape as follows: For $A_{s}$, letting the number of sites on a side the scale $l$, the perimeter $P_{B}=4*l$. For $A_{r}$, letting the number of sites on one of the sides franking the right angle the scale $l$, $P_{B}=(2+\sqrt{2})(l+0.5)$. For $A_{t}$, letting the number of sites on the side corresponding to its height the scale $l$, $P_{B}=(4l-1)+\sqrt{2}l=(4+\sqrt{2})l-1$. For $A_{p}$, letting the number of sites on a side in the direction of the base the scale $l$, $P_{B}=2(1+\sqrt{2})l$ For $A_{s}$ and $A_{r}$, we take as the system $A$ a square with the 3 times larger than respective scale, while for $A_{t}$ and $A_{p}$, a rectangle of 4 times scale in the direction of the base and 3 times scale in the direction of the height. The sides of the squares and rectangles are assumed to be parallel to those of the universal system. The MIs and LNs are computed by varying the scale of the $A_{2}$ system from $l_{\text{min}}=6$ to $l_{\text{max}}=32$ for $A_{s}$ and $A_{r}$, and from $l_{\text{min}}=6$ to $l_{\text{max}}=27$ for $A_{t}$ and $A_{p}$. Fig.1 shows the largest $A_{2}$ systems and corresponding $A$ systems.
@@
We fit the values of MIs and LNs in the range of $P_{B}$ from the suitable minimum perimeter $P_{\text{min}}$ to the $P_{\text{end}}$ to Eqs (31) and (32) by the least squares method in order to obtain the constants $a,\tilde{a}$, $b_{\text{total}},\tilde{b}_{\text{total}}$, $c_{-k},\tilde{c}_{-k}$, $k=0,1,\ldots n$. We perform these operations changing $P_{\text{end}}$ up to $P_{\text{max}}$, the perimeter corresponding to the largest scale $l_{\text{max}}$ of the subsystem $A_{2}$ and obtain the mean values and standard deviations(STDs) of $\tilde{b}_{\text{total}}$, $b_{\text{total}}$. In the range from 0 to 4 of $n$, the smallest STD of $\tilde{b}_{\text{total}}$s of (31) which determine how far $k$ of which parameters $\tilde{c}_{-k}$ are to be fitted, for all shapes of $A_{2}$ are obtained for $n=1$. On the other hand, those of $b_{\text{total}}$s of (32) are $n=2$ for $A_{s}$ and $A_{r}$, $n=1$ for $A_{t}$ and $A_{p}$. Hereafter, the values of $\tilde{b}_{\text{total}},b_{\text{total}}$ and their STDs are denoted by $\tilde{b}_{s},b_{s}$, $\tilde{b}_{ss},b_{ss}$, $\tilde{b}_{r},b_{r}$, $\tilde{b}_{rs},b_{rs}$, etc., respectively, depending on the shape of $A_{2}$.
Let us discuss the corner function values obtained from LNs, which have smaller STDs, i.e., have more stable values than those obtained from MIs, which we discuss later. Corner functions $b(\pi/2)$, $b(\pi/4)$ and their STDs are given as follows.
$\begin{split}b(\pi/2)=\ &b_{s}/4,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ b(\pi/2)_{\text{std}}=b_{ss}/4,\\ b(\pi/4)=\ &b_{r}/2-b_{s}/8,\ \ \ \ \ \ \ \ \ b(\pi/4)_{\text{std}}=\sqrt{(b_{rs}/2)^{2}+(b_{ss}/8)^{2}},\end{split}$ (33)
while $b(3\pi/4)$ and $b(3\pi/4)_{\text{std}}$ can be given in two ways using $b_{t}$, $b_{ts}$ or $b_{p}$$b_{ps}$:
$\begin{split}b(3\pi/4)=\ &b_{t}-b_{r}/2-3b_{s}/8,\ \ b(3\pi/4)_{\text{std}}=\sqrt{b_{ts}^{2}+(b_{rs}/2)^{2}+(3b_{ss}/8)^{2}},\\ b(3\pi/4)=\ &b_{p}/2-b_{r}/2+b_{s}/8,\ \ \ \ b(3\pi/4)_{\text{std}}=\sqrt{(b_{ps}/2)^{2}+(b_{rs}/2)^{2}+(b_{ss}/8)^{2}},\end{split}$ (34)
@@
$\begin{split}b(\pi/2)=0.02954,&\ \ \ b(\pi/2)_{\text{std}}=0.00013,\\ b(\pi/4)=0.09670,&\ \ \ b(\pi/4)_{\text{std}}=0.00032,\end{split}$ (35)
while the two values of (34) are respectively given by
$\begin{split}b(3\pi/4)=0.00604,&\ \ \ \ b(3\pi/4)_{\text{std}}=0.00047,\\ b(3\pi/4)=0.00635,&\ \ \ \ b(3\pi/4)_{\text{std}}=0.00038.\end{split}$ (36)
These corner function values are almost equal to those in the Ref.[11], and are stable as the relative standard deviations are less than 0.004 or less than 0.07. However, in detail, the difference of $b(3\pi/4)$ from the literature is less than its STD, while those of $b(\pi/2)$ and $b(\pi/4)$ are about 4 and 3 times larger than their STDs, respectively.
Let us now discuss the corner functions obtained from Renyi entropy-based MIs. In particular, the corner functions $\tilde{b}(\theta)$ due to Renyi entropy with the order $1/2$ coincide with $b(\theta)$ due to LNs so that $b(\theta)$ and $\tilde{b}(\theta)$ are reasonable checks for each other [11]. Since for those corner functions $\tilde{b}(2\pi-\theta)=\tilde{b}(\theta)$, $0\leq\theta\leq\pi$ and $\tilde{b}_{\text{total}}$ of the Eq.(31) are rewritten to
@@
Then, the relationship between the values of each $\tilde{b}(\theta),\theta=\pi/2,\pi/4,3\pi/4$ and $\tilde{b}_{\text{total}}$s are obtained by not only changing each ’$b$’ to $\tilde{b}$’ but also replacing each $b_{\text{total}}$ to $\tilde{b}_{\text{total}}/2$ in equations (33) and (34): $\tilde{b}(\pi/2)=\tilde{b}_{s}/8$, $\tilde{b}(\pi/2)_{\text{std}}=\tilde{b}_{ss}/8$, etc.. The corner function values obtained from MI in the above manner are as follows:
$\begin{split}\tilde{b}(\pi/2)=0.02961,&\ \ \ \tilde{b}(\pi/2)_{\text{std}}=0.00003,\\ \tilde{b}(\pi/4)=0.1004,&\ \ \ \tilde{b}(\pi/4)_{\text{std}}=0.0011,\end{split}$ (37)

### v2

@@ Abstract
We study the entanglement indices such as logarithmic negativities (LNs) and mutual informations (MIs) between two adjacent subsets in an isolated universal set $U$ of harmonic oscillators arranged on a two dimensional lattice within a sufficiently large square. First, we verify the values of the corner functions of angle $\pi/2,\pi/4,3\pi/4$ presented in the previous study which adopts periodic boundary conditions (PBCs) for the $U$. The values of each corner function obtained from LNs are nearly equal to those in the previous ones, while those of $3\pi/4$ from MIs are not sufficiently consistent with those computed from LNs. Next, for the case where the universal system $U$ satisfies the fixed boundary conditions(FBCs), we calculate LNs, MIs at several locations in $U$, compare them, especially corner functions with the values obtained in the PBCs case, and examine the effect of the fixed ends. In addition, we examine Renyi entropies for sets of three dimensional lattice sites, the corner functions and the edge terms with solid angles and dihedral angles $\pi/2,\pi/4$, respectively.
[…]
A60,A61
@@ 1 Introduction
Other quantity that indicates the degree of entanglement between A and its environment B is Renyi entropyRenyi , which is an extension of the entanglement entropy and is expressed by:
@@
In the case that two subsystems $A_{1}$ and $A_{2}$ are interior of the universal system $U$: $U\supset A=A_{1}+A_{2}$, there are two representative measures of entanglement between $A_{1}$ and $A_{2}$:logarithmic negativity (LN) and mutual information(MI)Eisert ; Groisman . The LN is given by
@@
After the seminal work by Hawking and BekensteinHawking ; Bekenstein , which showed that the entropy of a black hole is proportional to the area of its event horizon, the so-called holographic principle was proposed which asserted that quantum theoretical processes in $d_{U}$ dimensional space could be explained by processes in one smaller dimension,i.e.,the surface areatHooft ; Maldacena . Furthermore, in the two dimensional space($d_{U}=2$), when the boundary between two adjacent subsystems contains sharp corners of opening angles $\theta_{k},k=1,2,\ldots$, it is said that the many quantities related to the entanglement can be expanded as
$\alpha_{B}P_{B}/\delta-\sum_{k}\beta(\theta_{k})\log P_{B}/\delta+\cdots,$ (5)
@@
where $P_{B}$ is the the boundary length between $A$ and $U$ or between $A_{1}$ and $A_{2}$, $\delta$ is the UV cutoffBueno . The coefficient $\alpha_{B}$ correspondent to the area law depends on the cut-off $\delta$. On the other hand, the corner functions $\beta(\theta_{k})$s are universal, independent of the way the continuous limit to the underlying theory is taken. In three dimensional space $d_{U}=3$, when the boundary surface with area $P_{B}$ has edges with length $E_{j}$ and dihedral angles $\theta_{j},(j=1,2,\ldots)$ as well as sharp vertices with solid angles $\Omega_{k}$, the expansion is given by
$\alpha_{B}P_{B}/\delta^{2}+\sum_{k}\beta(\Omega_{k})\log P_{B}/\delta^{2}-\sum_{j}e(\theta_{j})E_{j}/\delta+\cdots,$ (6)
where the coner functions $\beta(\Omega_{k})$s are universal similar to the those of formula(5), while the edge term coefficients $e(\theta_{j})$s are notDevakul .
@@
For a two dimensional(2D) harmonic lattice with periodic boundary conditions (PBCs), which is a discretized massless scalar field in a torus-like space, the RefNobili shows that the values of the corner functions of $\theta=\pi/2,\pi/4,3\pi/4$ obtained from LNs are highly consistent with those of conformal field theory describing the scalar field. It is also claimed that the values of the corner functions obtained from MIs corresponding to the Renyi entropies of order 1/2 are consistent with the above LN-derived values. In the computing process, we take the relation between the scale $L$ of the universal set $U$, the scale $l$ of $A$ and the lattice spacing $a$ to be $L\gg l\gg a$, so that the calculation on the lattice can approximate the based field theory which is the continuous limit. With those scale conditions, even when the canonical variables on the lattice sites of $U$ satisfy the fixed boundary conditions(FBCs), the values of those corner function in the region sufficiently far from the boundary of $U$ are expected to be consistent with the above ones calculated under the PBCs.
@@
In this paper, according to Ref.Nobili , we first consider, as $A_{2}$, a set of lattice sites in several figures with corners of open angles $\theta=\pi/2,\pi/4,3\pi/4$ on their boundaries and scales defined by the number of sites on the shortest side ranging from six to a maximum of about thirty. We suppose that $A$ is a set of sites in a square whose scale is 3 to 4 times larger than that of $A_{2}$ and sites of the set $A_{1}=A-A_{2}$ encloses $A_{2}$ sites. We calculate LNs and MIs between $A_{1}$ and $A_{2}$ for each figure while changing the scales of $A_{2}$ and A. We fit those quantities by the formula (5) to obtain the corner functions and their standard deviations in order to verify those values in the reference. Next, for the case where the canonical variables on the lattice sites satisfy the FBCs, we perform the same procedure as before at many locations in $U$ and observe whether the values of each corner function near the center of $U$ are consistent with those under the PBCs and how they change approaching the boundary of $U$.
@@
This paper is organized as follows: In section 2, we take the Hamiltonian of harmonic lattice which is the discretized massless scalar field. We present analytical solutions $p_{\vec{r}}$ and $q_{\vec{r}}$ of the Hamilton equations under PBCs and FBCs, as well as their vacuum correlation functions $X=\braket{0|q_{\vec{r}},q_{\vec{s}}|0}$ and $P=\braket{0|p_{\vec{r}},p_{\vec{s}}|0}$, where $\vec{r}$ and $\vec{s}$ are position vectors of lattice sites. Given the specific shapes and scales of the universal set $U$ and subsets $A$, $A_{1}$, $A_{2}$, after reviewing the prescription to compute LN and various entropies from $X$ and $P$ matrices, we obtain the value of each corner function of $\theta=\pi/2,\pi/4,3\pi/4$ (and each standard deviation) for harmonic lattice with PBCs and examine whether they are consistent with the previous study. In section 3, we compute the corner functions at many locations in $U$ which is a set of lattice sites under FBCs by the same computational prescription in Section 2, and analyze them. In section 4, we examine, in a 3D lattice, the corner functions and the edge term coefficients, respectively, of solid angles and of dihedral angles $\pi/2$$\pi/4$ of Renyi entropy of order $1/2$. In section 5, we present some conclusions.
@@ 2 Models of harmonic lattices and entanglement indices computation
and correlation functions of generalized coordinate $q_{\vec{r}}$ and their conjugate momentum $p_{\vec{r}}$ expressed as followsNobili :
@@
For the ground state of (21), the values of correlators (14),(15), (19),(20) are also to be obtained from the matrix $K$ as follows Lykken ; Yazdi :
@@
Consider a set of sites inside the universal system $U$ described in the previous section, and denote it as a subsystem $A$. Also, let its environment B=UAB=U-A. Using the correlators (14),(15), (19),(20), or (23),(24), we can obtain the Renyi entropies $S_{A}^{(n)}$and entanglement entropy S(A)S(A)Bombelli ; Nez . First, we make the matrices of rows and columns of the above correlators restricted to those corresponding to the sites of the subsystem, i.e., $X_{\vec{u}\vec{v}}$ and $P_{\vec{u}\vec{v}}$, where $\vec{u},\vec{v}$ are the positions of sites in the subsystem $A$. Second, we compute the matrix $\Xi=\sqrt{XP}$ and its eigenvalues $\xi_{i}$. Using the eigenvalues, we can describe the entanglement entropy(2) and Renyi entropies(3) as follows:
@@
$S_{A}^{(n)}$ $=\frac{1}{n-1}\sum_{i=1}^{\dim(\mathcal{H}_{A})}\log\left((\xi_{i}+\frac{1}{2})^{n}-(\xi_{i}-\frac{1}{2})^{n}\right)$ (26)
Furthermore, we divide the subsystem $A$ into subsubsystems $A_{1}$ and $_{2}, A=A$$A_{1}+A_{2}$. Let us consider mutual information (MI) and logarithmic negativity (LN) as indices of the entanglement between $A_{1}$ and $A_{2}$Groisman ; Nobili . The MI and corresponding combination of Renyi entropies are given by
@@
From the matrices $X,P$, we can also calculate LN between the subsystems. First, we find the partial transposition of the density matrix $\rho_{A}$ with respect to $A_{2}$, which corresponds to the time reversal only to the degrees of freedom in $A_{2}$Simon . Therefore, it does not change the matrix $X$. On the otherhand, as for the matrix $P$, when we rewrite the matrix into the form as:
@@
where $\tilde{b}(\theta_{i}^{(j)})$ and $b(\theta_{i}^{(j)})$ are so called corner functions and $\theta^{(j)}_{i}$ is the angle corresponding to the $i$-th vertex of $SB$ belongs to $A_{j}$Nobili .
@@
We define the scale $l$ and its perimeter $P_{B}$ of each shape as follows: For $A_{s}$, letting the number of sites on a side the scale $l$, the perimeter $P_{B}=4*l$. For $A_{r}$, letting the number of sites on one of the sides franking the right angle the scale $l$, $P_{B}=(2+\sqrt{2})(l+0.5)$. For $A_{t}$, letting the number of sites on the side corresponding to its height the scale $l$, $P_{B}=(4l-1)+\sqrt{2}l=(4+\sqrt{2})l-1$. For $A_{p}$, letting the number of sites on a side in the direction of the base the scale $l$, $P_{B}=2(1+\sqrt{2})l$. For $A_{s}$ and $A_{r}$, we take as the system $A$ a square with the 3 times larger than respective scale, while for $A_{t}$ and $A_{p}$, a rectangle of 4 times scale in the direction of the base and 3 times scale in the direction of the height. The sides of the squares and rectangles are assumed to be parallel to those of the universal system. The MIs and LNs are computed by varying the scale of the $A_{2}$ system from $l_{\text{min}}=6$ to $l_{\text{max}}=32$ for $A_{s}$ and $A_{r}$, and from $l_{\text{min}}=6$ to $l_{\text{max}}=27$ for $A_{t}$ and $A_{p}$. Fig.1 shows the largest $A_{2}$ systems and corresponding $A$ systems.
@@
We fit the values of MIs and LNs in the range of $P_{B}$ from the suitable minimum perimeter $P_{\text{min}}$ to the $P_{\text{end}}$ to Eqs (31) and (32) by the least squares method in order to obtain the constants $a,\tilde{a}$, $b_{\text{total}},\tilde{b}_{\text{total}}$, $c_{-k},\tilde{c}_{-k}$, $k=0,1,\ldots n$. We perform these operations changing $P_{\text{end}}$ up to $P_{\text{max}}$, the perimeter corresponding to the largest scale $l_{\text{max}}$ of the subsystem $A_{2}$ and obtain the mean values and standard deviations(SDs) of $\tilde{b}_{\text{total}}$, $b_{\text{total}}$. In the range from 0 to 4 of $n$, the smallest SD of $\tilde{b}_{\text{total}}$s of (31) which determine how far $k$ of which parameters $\tilde{c}_{-k}$ are to be fitted, for all shapes of $A_{2}$ are obtained for $n=1$. On the other hand, those of $b_{\text{total}}$s of (32) are $n=2$ for $A_{s}$ and $A_{r}$, $n=1$ for $A_{t}$ and $A_{p}$. Hereafter, the values of $\tilde{b}_{\text{total}},b_{\text{total}}$ and their SDs are denoted by $\tilde{b}_{s},b_{s}$, $\tilde{b}_{ss},b_{ss}$, $\tilde{b}_{r},b_{r}$, $\tilde{b}_{rs},b_{rs}$, etc., respectively, depending on the shape of $A_{2}$.
Let us discuss the corner function values obtained from LNs, which, as will be discussed later, have smaller SDs, i.e., have more stable values than those obtained from MIs. Corner functions $b(\pi/2)$, $b(\pi/4)$ and their SDs are given as follows.
$\begin{split}b(\pi/2)=\ &b_{s}/4,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ b(\pi/2)_{\text{sd}}=b_{ss}/4,\\ b(\pi/4)=\ &b_{r}/2-b_{s}/8,\ \ \ \ \ \ \ \ \ b(\pi/4)_{\text{sd}}=\sqrt{(b_{rs}/2)^{2}+(b_{ss}/8)^{2}},\end{split}$ (33)
while $b(3\pi/4)$ and $b(3\pi/4)_{\text{sd}}$ can be given in two ways using $b_{t}$, $b_{ts}$ or $b_{p}$$b_{ps}$:
$\begin{split}b(3\pi/4)=\ &b_{t}-b_{r}/2-3b_{s}/8,\ \ b(3\pi/4)_{\text{sd}}=\sqrt{b_{ts}^{2}+(b_{rs}/2)^{2}+(3b_{ss}/8)^{2}},\\ b(3\pi/4)=\ &b_{p}/2-b_{r}/2+b_{s}/8,\ \ \ \ b(3\pi/4)_{\text{sd}}=\sqrt{(b_{ps}/2)^{2}+(b_{rs}/2)^{2}+(b_{ss}/8)^{2}},\end{split}$ (34)
@@
$\begin{split}b(\pi/2)=0.02954,&\ \ \ b(\pi/2)_{\text{sd}}=0.00013,\\ b(\pi/4)=0.09670,&\ \ \ b(\pi/4)_{\text{sd}}=0.00032,\end{split}$ (35)
while the values of (34) are respectively given by
$\begin{split}b(3\pi/4)=0.00604,&\ \ \ \ b(3\pi/4)_{\text{sd}}=0.00047,\\ b(3\pi/4)=0.00635,&\ \ \ \ b(3\pi/4)_{\text{sd}}=0.00038.\end{split}$ (36)
These corner function values are almost equal to those in the Ref.Nobili , and are stable as the relative standard deviations are less than 0.004 or less than 0.07. However, in detail, the difference of $b(3\pi/4)$ from the literature is less than its SD, while those of $b(\pi/2)$ and $b(\pi/4)$ are about 4 and 3 times larger than their SDs, respectively.
Let us now discuss the corner functions obtained from Renyi entropy-based MIs. In particular, the corner functions $\tilde{b}(\theta)$ due to Renyi entropy with the order $1/2$ should coincide with $b(\theta)$ due to LNs so that $b(\theta)$ and $\tilde{b}(\theta)$ are reasonable checks for each other Nobili . Since for those corner functions $\tilde{b}(2\pi-\theta)=\tilde{b}(\theta)$, $0\leq\theta\leq\pi$ and $\tilde{b}_{\text{total}}$ of the Eq.(31) are rewritten to
@@
Then, the relationship between the values of each $\tilde{b}(\theta),\theta=\pi/2,\pi/4,3\pi/4$ and $\tilde{b}_{\text{total}}$s are obtained by not only changing each ’$b$’ to $\tilde{b}$’ but also replacing each $b_{\text{total}}$ to $\tilde{b}_{\text{total}}/2$ in equations (33) and (34): $\tilde{b}(\pi/2)=\tilde{b}_{s}/8$, $\tilde{b}(\pi/2)_{\text{sd}}=\tilde{b}_{ss}/8$, etc.. The corner function values obtained from MI in the above manner are as follows:
$\begin{split}\tilde{b}(\pi/2)=0.02962,&\ \ \ \tilde{b}(\pi/2)_{\text{sd}}=0.00003,\\ \tilde{b}(\pi/4)=0.1004,&\ \ \ \tilde{b}(\pi/4)_{\text{sd}}=0.0011,\end{split}$ (37)